# Fast Revenue Maximization

### v1

@@ 2.2 Reduction to 2D-problem
Consider $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and consider the extended information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}:=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup\{({r^{*}},{q^{*}})\}$. We note that $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ is the subset of $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ where distributions also satisfy $\overline{F}({r^{*}})={q^{*}}$. We also let $\mathscr{A}=\{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\text{ s.t. }{r^{*}}\in\mathscr{O}(F)\text{ and }\overline{F}({r^{*}})={q^{*}}\}.$ We first show that,
$\inf_{F\in\mathscr{A}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]\geq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)].$ (2.8)
@@
$\inf_{F\in\mathscr{A}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]\leq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)].$ (2.9)
@@
We first show that $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ is non-empty. Note that since $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ there exists $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that $\overline{F}({r^{*}})={q^{*}}$. Such a distribution belongs to $\mathscr{A}$ and therefore belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$. Therefore, this set is not empty. It then follows from Lemma 3 that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
$\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)=\max\left(\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i},{r^{*}}\cdot{q^{*}}\right).$
Furthermore, we argue that ${r^{*}}\cdot{q^{*}}\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$. This inequality holds because $\mathscr{A}$ is non-empty and for any distribution $F\in\mathscr{A}$, we have for every $i\in\{0,\ldots,N+1\}$ that ${r^{*}}\cdot{q^{*}}={r^{*}}\cdot\overline{F}({r^{*}})\stackrel{{\scriptstyle(a)}}{{\geq}}p_{i}\cdot\overline{F}(p_{i})=p_{i}\cdot q_{i},$ where $(a)$ holds because $F\in\mathscr{A}$ and hence ${r^{*}}$ is an optimal price for this distribution. This implies that $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)={r^{*}}\cdot{q^{*}}$ and given that $\overline{L}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})={q^{*}}$, it concludes that ${r^{*}}$ is an optimal price for $\overline{L}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ with associated conversion rate ${q^{*}}$.
@@ 2.3 Reduction to a 1D problem
$\rho:\begin{cases}[0,1]\to\mathbb{R}_{+}\\ {q^{*}}\mapsto\mathbb{E}_{p\sim\Psi}\left[\lambda\cdot{r^{*}}\cdot{q^{*}}-\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right].\end{cases}$
@@
First, we remark from the epigraph formulation in Appendix A that $\frac{\rho(q^{*})}{q^{*}}$ is non-negative. To prove that $\frac{\rho(q^{*})}{q^{*}}$ is non-decreasing remark that, $\frac{\rho({q^{*}})}{{q^{*}}}=\lambda r^{*}-\mathbb{E}_{p\sim\Psi}\left[\frac{1}{q^{*}}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right]$. Hence, it is sufficient to show that for every $p$, the mapping $\tilde{\rho}:{q^{*}}\mapsto\frac{1}{{q^{*}}}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)$ is non-increasing.
When $\pazocal{C}=\pazocal{G}$, $\tilde{\rho}$ is a constant when $\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))={q^{*}}$ and is the inverse function when $\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))=q_{i}\neq{q^{*}}$. In both cases, it is non-increasing.
@@
$\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)=p\cdot\overline{G}_{\pazocal{C}}(p|(\hat{p}_{i},\hat{q}_{i}),(\hat{p}_{i+1},\hat{q}_{i+1})).$
Hence, $\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)$ only varies as a function of ${q^{*}}$ if ${r^{*}}\in\{\hat{p}_{i},\hat{p}_{i+1}\}$. In all other cases, the numerator in $\tilde{\rho}$ is a constant which does not depend on ${q^{*}}$ and therefore $\tilde{\rho}$ is clearly non-increasing. We next show that when either $p_{i}={r^{*}}$ or $p_{i+1}={r^{*}}$, $\tilde{\rho}$ is also non-increasing.
@@
Case 2: $p_{i+1}={r^{*}}$. Using again one of the alternative forms of $\bar{G}_{\pazocal{C}}$ derived in the proof of Lemma 1 we obtain that,
@@
$\sup_{F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{r^{*}\in[\underline{v},\overline{v}]}\sup_{q^{*}\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\rho(q^{*})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sup_{r^{*}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot\rho(\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}}))$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sup_{r^{*}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot\rho(\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})).$
@@
$\overline{F}_{\pazocal{C}}(v|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right).$ (2.10)
@@
$\sup_{F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{{r^{*}}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=\sup_{{r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),$
where the last equality holds by defining $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\{{r^{*}}\in[\underline{v},\overline{v}]\text{ s.t. }{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}$. This concludes our reduction to a one-dimensional optimization problem.
@@
For every ${r^{*}}\in[\underline{v},\overline{v}]$, the ${r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ if and only if,
$i)$ $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})))$ is non-empty, and $ii)$ ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$.
@@
First, if ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$, Lemma 7 implies that $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. By definition, of ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ this implies that $i)$ and $ii)$ are satisfied.
@@
Reciprocally, $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})))$ is non-empty, Lemma 3 implies that the distribution $\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right)\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})))\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Furthermore, the fact Lemma 4 along with the condition that ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$, implies that ${r^{*}}$ is an optimal price for $\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right)$ with associated conversion-rate $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. This proves that $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Hence, this set is non-empty. ∎

### v2

@@ 2.2 Reduction to a two-dimensional problem
Consider $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and consider the extended information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}:=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup\{({r^{*}},{q^{*}})\}$. We note that $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ is the subset of $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ where distributions must also satisfy $\overline{F}({r^{*}}-)={q^{*}}$. We also let $\mathscr{A}=\{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\text{ s.t. }{r^{*}}\in\mathscr{O}(F)\text{ and }\overline{F}({r^{*}}-)={q^{*}}\}.$ We first show that,
$\inf_{F\in\mathscr{A}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]\geq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)].$ (2.6)
@@
$\inf_{F\in\mathscr{A}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]\leq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)].$ (2.7)
@@
We first show that $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ is non-empty. Note that since $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ there exists $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that $\overline{F}({r^{*}}-)={q^{*}}$. Such a distribution belongs to $\mathscr{A}$ and therefore belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$. Therefore, this set is not empty. It then follows from Lemma 3 that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
$\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)=\max\left(\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i},{r^{*}}\cdot{q^{*}}\right).$
Furthermore, we argue that ${r^{*}}\cdot{q^{*}}\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$. This inequality holds because $\mathscr{A}$ is non-empty and for any distribution $F\in\mathscr{A}$, we have for every $i\in\{0,\ldots,N+1\}$ that ${r^{*}}\cdot{q^{*}}={r^{*}}\cdot\overline{F}({r^{*}}-)\stackrel{{\scriptstyle(a)}}{{\geq}}p_{i}\cdot\overline{F}(p_{i}-)=p_{i}\cdot q_{i},$ where $(a)$ holds because $F\in\mathscr{A}$ and hence ${r^{*}}$ is an optimal price for this distribution. This implies that $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)={r^{*}}\cdot{q^{*}}$ and given that $\overline{L}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})={q^{*}}$, it concludes that ${r^{*}}$ is an optimal price for $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ with associated conversion rate ${q^{*}}$.
@@ 2.3 Reduction to a one-dimensional problem
$\rho:\begin{cases}[0,1]\to\mathbb{R}_{+}\\ {q^{*}}\mapsto\mathbb{E}_{p\sim\Psi}\left[\lambda\cdot{r^{*}}\cdot{q^{*}}-\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right].\end{cases}$
@@
First, we remark from the epigraph formulation in Section 3 that $\frac{\rho(q^{*})}{q^{*}}$ is non-negative. To prove that $\frac{\rho(q^{*})}{q^{*}}$ is non-decreasing remark that, $\frac{\rho({q^{*}})}{{q^{*}}}=\lambda r^{*}-\mathbb{E}_{p\sim\Psi}\left[\frac{1}{q^{*}}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right]$. Hence, it is sufficient to show that for every $p$, the mapping $\tilde{\rho}:{q^{*}}\mapsto\frac{1}{{q^{*}}}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)$ is non-increasing.
When $\pazocal{C}=\pazocal{G}$, $\tilde{\rho}$ is a constant when for every $p$ such that $\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))={q^{*}}$ and is the inverse function when $\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))=q_{i}\neq{q^{*}}$. In both cases, it is non-increasing.
@@
$\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)=p\cdot\overline{G}_{\pazocal{C}}(p-|(\hat{p}_{i},\hat{q}_{i}),(\hat{p}_{i+1},\hat{q}_{i+1})).$
Hence, $\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)$ only varies as a function of ${q^{*}}$ if ${r^{*}}\in\{\hat{p}_{i},\hat{p}_{i+1}\}$. In all other cases, the numerator in $\tilde{\rho}$ is a constant which does not depend on ${q^{*}}$ and therefore $\tilde{\rho}$ is clearly non-increasing. We next show that when either $p_{i}={r^{*}}$ or $p_{i+1}={r^{*}}$, $\tilde{\rho}$ is also non-increasing.
@@
Case 2: $p_{i+1}={r^{*}}$. Using again one of the alternative forms of $\overline{G}_{\pazocal{C}}$ derived in the proof of Lemma 1 we obtain that,
@@
$\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{r^{*}\in[\underline{v},\overline{v}]}\sup_{q^{*}\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\rho(q^{*})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sup_{r^{*}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot\rho(\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}}))$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sup_{r^{*}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot\rho(\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})).$
@@
$\overline{F}_{\pazocal{C}}(v|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right).$ (2.8)
@@
$\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{{r^{*}}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=\sup_{{r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),$
where the last equality holds by defining $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\{{r^{*}}\in[\underline{v},\overline{v}]\text{ s.t. }{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}$. This concludes our reduction to a one-dimensional optimization problem and concludes the proof of Theorem 1.
@@
Fix ${r^{*}}\in[\underline{v},\overline{v}]$. Then ${r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ if and only if,
$i)$ $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})))$ is non-empty, and $ii)$ ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$.
@@
First, if ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$, Lemma 7 (in Section 6) implies that $\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. By definition, of ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ this implies that $i)$ and $ii)$ are satisfied.
@@
Reciprocally, if $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})))$ is non-empty, then Lemma 3 implies that the distribution $\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right)\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})))\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Furthermore, Lemma 4 along with the condition that ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$, implies that ${r^{*}}$ is an optimal price for $\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right)$ with associated conversion-rate $\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. This proves that $\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Hence, this set is non-empty. ∎
3 Quantifying the Value of Information
Theorem 1 implies that one can efficiently quantify the worst-case performance of any pricing mechanism given a particular information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. We next, establish how this allows to derive a lower bound on (Maximin). We note that (Maximin) characterizes the intrinsic value of an information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ as this ratio represents the best achievable performance across all pricing mechanisms having access to past prices and conversion rates in $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. Evaluating (Maximin) will be critical to derive new insights on price experimentation in Section 4.
The following discussion presents an application of the method developed in (Allouah et al., 2023) to our setting. First, we note that the epigraph formulation of (Maximin) satisfies,
$\!\sup_{\Psi(\cdot)\in\pazocal{P},\,\lambda\in[0,1]}$ $\lambda$
$\!\sup_{\Psi(\cdot)\in\pazocal{P},\,\lambda\in[0,1]}$ $\lambda$
subject to $R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\leq 0\quad\text{ for every }{r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}}).$
Finally, we derive a lower bound on (Maximin) by restricting the space of randomized mechanisms to discrete mechanisms (i.e., mechanisms such that the distribution of prices is discrete over $[\underline{v},\overline{v}]$) and by projecting the one dimensional set of constraints into the associated support of the mechanisms. More specifically, consider an increasing sequence of $M+2$ positive reals $\mathbb{A}=\{a_{i}\}_{i=0}^{M+1}$ forming a partition of the feasible interval $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ for some $M\in\mathbb{N}^{*}$ and define the set of discrete mechanisms on $\mathbb{A}$ as
$\pazocal{P}_{\mathbb{A}}=\left\{\Psi\in\pazocal{P}\>:\>\Psi(x)=\sum_{i=0}^{M+1}\psi_{i}\cdot\mathbf{1}\{x\geq a_{i}\},\mbox{ for some }0\leq\psi_{i}\leq 1,\>\sum_{j=0}^{M+1}\psi_{i}=1\right\}.$
$\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})=\;$ $\!\sup_{\mathbf{p},\lambda}$ $\lambda$ (3.3a)
s.t. $\lambda\cdot\mbox{opt}\left(F_{\pazocal{C}}(\cdot|a_{i+1}^{-},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)-\sum_{j=0}^{M}\mathrm{Rev}\left(a_{j}|F_{\pazocal{C}}(\cdot|a_{i},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\psi_{j}\leq 0\quad i\in\{0,\ldots,M+1\},$ (3.3b)
$\sum_{j=0}^{M+1}\psi_{j}\leq 1,\quad\psi_{i}\geq 0\quad i\in\{0,\ldots,M+1\},\quad 0\leq\lambda\leq 1,$ (3.3c)
where $\mbox{opt}(F_{\pazocal{C}}(\cdot|a_{i+1}^{-},\mathscr{I}_{\mathbf{p},\mathbf{q}}))=\lim_{x\to a_{i+1}^{-}}\mbox{opt}(F_{\pazocal{C}}(\cdot|x,\mathscr{I}_{\mathbf{p},\mathbf{q}}))$ for any $i=0,...M$.
We note that a line of technical arguments similar to the one used in (Allouah et al., 2023) enables to show that $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})$ is a lower bound on the value of (Maximin). Furthermore, the solution $(\psi_{i})_{i\leq N+2}$ provides a feasible distribution of prices. Finally, we note that by judiciously constructing a discrete grid one can ensure that the values associated with the sequence of linear programs converge to the one of (Maximin) as the grid becomes finer.